13x+4x-5.2=53+12.5 What is the value of x?
Did I do this problem right?
13x+4x-5.2=53+12.5+5.2
17x=69.7
17/17 : 69.7/17
x=4.1

Answer by stanbon(75887)

13x+4x-5.2=53+12.5 What is the value of x?
Did I do this problem right?
13x+4x-5.2 = 53+12.5+5.2
==========================================---
It looks like you added 5.2 to the right side but not the left side.
---
13x+4x-5.2+5.2 = 53+12.5+5.2
----
17x = 70.7
x = 4.159...
=================
Cheers,
Stan H.